OpenStudy (anonymous):

Given ABCD ~ quadrilateral JKLM and AD =14, JM =6 and KL=9, what is BC?

OpenStudy (superdavesuper):

ABCD ~ JKLM so their sides are in the same ratio which means \[\frac{ AD }{ JM } = \frac{ BC }{ KL }\] insert the values n solve for BC

OpenStudy (superdavesuper):

u need to find BC...use my eqn above to solve plz

OpenStudy (superdavesuper):

\[\frac{ AD }{ JM } = \frac{ BC }{ KL }\] and AD =14, JM =6 and KL=9

OpenStudy (superdavesuper):

can u put the number into the eqn and solve for BC?

OpenStudy (anonymous):

21

OpenStudy (superdavesuper):

correct - 21 is the right ans :)